OpenStudy (anonymous):

Given the quadratic function f(x) = 25x2 – 100x + 100, find a value of x such that f(x) = 25.

OpenStudy (anonymous):

no

OpenStudy (zzr0ck3r):

what are you doing in class?

OpenStudy (anonymous):

You know factorization of a quadratic expression?

OpenStudy (anonymous):

maybe give me a min.

OpenStudy (anonymous):

my answer is not making any sense

OpenStudy (zzr0ck3r):

what are you doing in class? the "AC" method, the quadratic formula, or completing the square?

OpenStudy (anonymous):

x^2=-1

OpenStudy (zzr0ck3r):

please answer my question then I can help you

OpenStudy (anonymous):

quadratic formula

OpenStudy (zzr0ck3r):

ok we have (-b +- sqrt(b^2-4ac))/2a right?

OpenStudy (anonymous):

yep

OpenStudy (zzr0ck3r):

what is a? what is b? and what is c?

OpenStudy (anonymous):

25

OpenStudy (zzr0ck3r):

x^2 -4x+3

OpenStudy (zzr0ck3r):

this is the equation, ax^2+bx, + c

OpenStudy (zzr0ck3r):

25x^2-100x+100 = 25 first we want to set it equal to 0 so we subtract 25 from both sides 25x^2-100x+75 = 0 then I would devide everything by 25 to make life nice. x^2 -4x+3 = 0

OpenStudy (zzr0ck3r):

so we are at x^2-4x+3 = 0

OpenStudy (anonymous):

a=x^2, b=-4x , c=3

OpenStudy (zzr0ck3r):

we just want the coefficients

OpenStudy (zzr0ck3r):

a = 1, b = -4, c = 3

OpenStudy (anonymous):

oh

OpenStudy (zzr0ck3r):

now we have (-b+-sqrt(b^2-4ac))/2a = (-(-4)+-sqrt((-4)^2-4*1*3))/(2*1) = (16+-sqrt(16-12))/2 = (16+-sqrt(4))/2 = (16+-2)/2 at this point we have two equations (16+2)/2 and (16-2)/2 so what is the final answer?

OpenStudy (anonymous):

some correction: -(-4)=4, (4+-sqrt(4))/2

OpenStudy (anonymous):

it will be 2 correct answers

OpenStudy (anonymous):

you all have lost me im gettin x=3,x=1

OpenStudy (zzr0ck3r):

*now we have (-b+-sqrt(b^2-4ac))/2a = (-(-4)+-sqrt((-4)^2-4*1*3))/(2*1) = (4+-sqrt(16-12))/2 = (4+-sqrt(4))/2 = (4+-2)/2 at this point we have two equations (4+2)/2 and (4-2)/2 we have 3 and 1

OpenStudy (zzr0ck3r):

correct

OpenStudy (anonymous):

yes ok

OpenStudy (zzr0ck3r):

note there is a much nicer way of doing this x^2-4x+3 = 0 x^2-x-3x+3 = 0 x(x-1) -3(x-1) = 0 (x-1)(x-3) = 0 x-1 = 0 and x-3 = 0 so x = 1 and x = 3

OpenStudy (anonymous):

TO check your answers you can plug 1 and 3 into original eq-n and will receive 25